OpenStudy (christos):

Linear Algebra, Determine a constant 'k' such that (kA)^T * (kA) = 1 A = -2 1 -1 Are there more than one such constants ?

hartnn (hartnn):

[-2k k -k] [-2k ] [ k ] = 4k^2 +k^2 +k^2 = 1 [ -k ] got that?

OpenStudy (christos):

yes @hartnn

hartnn (hartnn):

so you'll get 2 values of k from there ..

OpenStudy (christos):

@hartnn is it true that if a homogeneous system has at least one more solution than the trivial one, then it has infinitely many solutions ?

hartnn (hartnn):

a homogeneous system, either has no solution or only trivial solution or infinitely many solutions

OpenStudy (christos):

could be non square and stil hve the same properties, right ?

hartnn (hartnn):

yes

OpenStudy (kittiwitti1):

Do you still need help on this problem?

rebeccaxhawaii (rebeccaxhawaii):

^^

OpenStudy (kainui):

I suggest not distributing k to the elements of the vector, \[(kA)^\top kA = 1\] \[k^2 A^\top A = 1\] \[k^2 = \frac{1}{A^\top A}\] \[k = \frac{\pm 1}{\sqrt{A^\top A}}\] Don't let the matrix notation confuse you, \(A^\top A \) is just the dot product of the vector A with itself, and the dot product is a scalar.

OpenStudy (christos):

thank you guys